INDUCED INFRARED SPECTRA OF DOUBLE VIBRATIONAL TRANSITIONS IN H_{2} − N_{2} AND D_{2} − N_{2}

\underline{C. STAMP}, R. D. G. PRASAD, P. G. GILLARD, and S. PADDI REDDY, Department of Physics and Physical Oceanography, Memorial University of Newfoundland, St. John’s, NF, A1B 3X7, Canada.

Collision-induced infrared absorption spectra of the double vibrational transitions of H_{2} (v=1 ← 0) + D_{2} (v=1 ← 0) have been investigated in their binary mixtures in the spectral region 5900 - 7100 cm^{−1}. The spectra were recorded with a 2 m absorption cell for partial densities of H_{2} and N_{2} in the range 60 to 315 and 100 to 315 amagat. respectively, at room temperature. The observed spectra are interpreted in terms of the following transitions: O_{1}(J) of H_{2}+Q_{1}(J) of N_{2}; Q_{1}(J) of H_{2}+O_{1}(J) of N_{2}; Q_{1}(J) of H_{2}+Q_{1}(J) of N_{2}; Q_{1}(J) of H_{2}+S_{1}(J) of N_{2}; S_{1}(J) of H_{2}+Q_{1}(J) of N_{2}.

Analysis of the absorption profiles has been carried out. Similar experiments with binary mixtures of D_{2} + N_{2} are also in progress. The results of this work including the absorption coefficients, line-shape function, etc., will be presented.